In a crank and slotted lever quick return motion mechanism, the distance between the fixed

centres O and C is 200 mm. The driving crank CP is 75 mm long. The pin Q on the slotted lever,

360 mm from the fulcrum O, is connected by a link QR 100 mm long, to a pin R on the ram. The

line of stroke of R is perpendicular to OC and intersects OC produced at a point 150 mm from C.

Determine the ratio of times taken on the cutting and return strokes.


1
Expert's answer
2021-05-12T07:51:33-0400

Quick return ratio = Time of cutting/time of return = "\dfrac{a}{b} = 360-\dfrac{a}{a}"


OC =200 mm ,CP = 75mm


"\cos \dfrac{b}{2} = \dfrac{CP}{OC} = \dfrac{75}{200}"


"\dfrac{b}{2}= 68°"


"b = 136\u00b0"


QRatio = "\dfrac{360-136}{136} = 1.647"